The market for candy is perfectly competitive, and the current market price of candy is $10. A particular firm has a short-run marginal cost of production of MC = 0.2q, where q is the number of bicycles produced by the firm.

a. If it is optimal for the firm to produce a positive amount of output in the short run, how much should it produce?

b. Suppose that the firm has fixed costs of $30, and its average variable cost when producing q candies is given by AVC = 0.1q (so, e.g., when q = 10 the firm has AVC = 1). Should the firm produce the amount that you found in part (a) or shut down (produce q = 0)?

c. Suppose instead that the firm has fixed costs of $50, and its variable cost when producing the number of candies found in part (a) is $200. What is the firm’s profit if it produces this amount of output? If this is a constant-cost industry and the demand curve does not change over time, will firms enter or exit this market in the long run?
Answer #1

Answer : a) For perfectly competitive firm at equilibrium P (Price) = MC occur. So,

10 = 0.2q

=> q = 10 / 0.2 = 50

Therefore, here the firm should produce 50 units of output.

b) The firm's short-run shutdown price is that price which is equal to AVC.

Firm's AVC = 0.1q = 0.1 * 50 = $5.

Here the price level is $10.

As here the market price level is higher than the firm's AVC hence the firm should produce 50 units of output level.

c) At 50 units of output level

Total cost = Fixed cost + Variable cost = 50 + 200 = $250.

Total revenue = Price * Quantity = 10 * 50 = $500.

Profit = Total revenue - Total cost = 500 - 250 = $250.

So, if the firm produce 50 units of output level then the firm's profit is $250.

If this insudtry has constant cost level and demand curve does not change then in long-run new firms will enter into the market. Because here in short-run the firm earns economoic profit.